Suppose your hamster poops X times before noon and Y times after noon. Where X is a Poisson (2) random variable and Y is an independent Poisson(3) random variable.

a. What is the expected number of poops the hamster took in a day?

b. What is the probability the hamster pooped fewer than 2 times a day?
Answer #1

a expected number of poops the hamster took in a day= E(X+Y)=E(X)+E(Y) =2+3 =5

b) probability the hamster pooped fewer than 2 times a day=P(X<2) =P(X=0)+P(X=1)

=e-5*500!+e-5*511! =0.0404